what point lies in the solution set for the following system:
4x+y<1
8y-3x>24

Answer by stanbon(48558)

4x+y<1
8y-3x>24
--------------
Rewrite as:
4x+y=1
-3x+8y=4
------------
Multiply 1st equation by 8 to get:
32x+8y=8
-3x+8y=4
---------
Subtract to get:
35x=4
x=4/35
--------
Substitute to solve for "y"
4(4/35)+y=1
y = 35/35-16/35 = 19/35
-----------
The lines cross at (4/35,19/35)
Sketch the line y=-4x+1 the point you want must be below it (less y value)
such as (4/35,0)
=========
Cheers,
Stan H.